A mass m is suspended from a wire. Change in length of the wire is dl. Now the same wire is stretched to double its length and the same mass is suspended from the wire. The change in length in this case will become :-


a)dl          b) 2dl                     c) 4dl                     d) 8dl

Grade:11

1 Answers

Chirrag Nangia
11 years ago

Let F = mg be the force

L , A° , dlº and 2L , A , dl be the length , area , chenge in length before and efter stretching

On stretching the wire the volume before and after will remain same

Thus A°*L=A*2L  ≡ A°=2A

Now equating youngs modulus as material of wire remain same

(F*L)/(A°*dl°) = (F*2L)/(A*dl)

2A°*dl° = A*dl

4dl° = dl